# Rotational dynamics of a disk in a thin film of weakly nematic fluid subject to linear friction

### v1

@@ V Exact solution in the absence of friction
Figure 3: Streamlines and quiver plots of the image flow field $\bm{v}^{*}$ for a disk rotating steadily in a two-dimensional, uniaxially anisotropic fluid. The results are shown for three values of the parameter $c=\alpha_{\parallel}a$, namely $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Additionally, two combinations of the parameters $A$ and $\beta$ quantifying the anisotropy of viscosity and friction are examined, namely $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Analytical solutions are depicted on the right-hand sides ($x>0$), while results from finite-element simulations are shown on the left-hand sides ($x<0$). We recall that $\bm{v}^{*}$ is rescaled by $a\Omega$ and that $c=\alpha_{\parallel}a$.
Figure 4: Streamlines and quiver plots of the image flow field $\bm{v}^{*}$ for a disk rotating steadily in a two-dimensional, uniaxially anisotropic fluid in the absence of friction, that is, for $\alpha_{\parallel}=\alpha_{\perp}=0$. Exact analytical solutions are shown on the right-hand sides ($x>0$), while results from finite-element simulations are shown on the left-hand sides ($x<0$). Again, $\bm{v}^{*}$ is rescaled by $a\Omega$.
@@ VI Comparison with results from finite-element simulations
In Fig. 3 we present streamlines and quiver plots of the image flow field for a disk undergoing steady rotational motion in a two-dimensional, uniaxially anisotropic fluid. The results are depicted for three values of the parameter $c=\alpha_{\parallel}a$, namely $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). We examine two combinations of the parameters quantifying the anisotropy of viscosity and friction, $A$ and $\beta$, namely $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Analytical solutions are displayed in the range $x>0$, while results from finite-element simulations are shown in the range $x<0$.
In Fig. 4, we present the image flow field for rotational motion in a nematic liquid crystal without friction, a scenario for which exact analytical expressions can be derived, see Sec. V. Results are illustrated for the two values of $A=4$ and $10$. They are sufficiently large for the perturbative solution not to remain quantitatively correct any more in general. The analytical solution, displayed on the right-hand sides, aligns perfectly with the finite-element simulations shown on the left-hand sides. Notably, the flow field features regions of large magnitudes of velocity. These regions are arranged along the edge of an octagon near the surface of the disk.
In the listed contexts, we concentrate on presenting the image flow field rather than the entire flow field. In this way, we enhance the visualization of the impact of the anisotropies. Our results reveal that the flow field is significantly influenced by changes in the parameters of anisotropy. In some cases, they exhibit vortical structures with closed streamlines and regions of high magnitude of velocity near the disk. Overall, our theoretical predictions based on a perturbative approach align very well with results from finite-element simulations that solve the complete set of hydrodynamic equations.
@@ VII Conclusions
During the further course of the topic and within a wider context, expansions of the theory to thin elastic membranes functionalized by rigid, plate-like inclusions represent a next step. We plan to extend our considerations to elastic composite systems in the future, particularly aiming at a theoretical consideration of membranes of globally aligned liquid crystal elastomers [73, 74, 75, 76, 77, 78].
@@
The Maple 24 scripts containing the final expressions for the velocity field are accessible through the Zenodo repository using the DOI XXX .

### v2

@@ V Exact solution in the absence of friction
Notably, Eq. (70) shows that $R=0$ when $A=-4$, which is the smallest value $A$ can take when requiring both physically and mathematically that $B\geq 0$. While providing a physical interpretation might be challenging, it is not surprising that the case $A=-4$ is special, given that it represents the lower bound of this interval of $A$.
@@
Figure 3: Streamlines and quiver plots of the leading-order correction to the flow field due to uniaxial anisotropy in viscosity and friction $\bm{v}^{*}$ for a disk rotating steadily in a two-dimensional, uniaxially anisotropic fluid. The results are shown for three values of the parameter $c=\alpha_{\parallel}a$, namely $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Additionally, two combinations of the parameters $A$ and $\beta$ quantifying the anisotropy of viscosity and friction are examined, namely $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Perturbative solutions are depicted on the right-hand sides ($x>0$), while results from finite-element simulations are shown on the left-hand sides ($x<0$). We recall that $\bm{v}^{*}$ is rescaled by $a\Omega$ and that $c=\alpha_{\parallel}a$.
Figure 4: Streamlines and quiver plots of the leading-order correction to the flow field due to uniaxial viscosity anisotropy $\bm{v}^{*}$ for a disk rotating steadily in a two-dimensional, uniaxially anisotropic fluid in the absence of friction, that is, for $\alpha_{\parallel}=\alpha_{\perp}=0$. Exact analytical solutions are shown on the right-hand sides ($x>0$), while results from finite-element simulations are shown on the left-hand sides ($x<0$). Again, $\bm{v}^{*}$ is rescaled by $a\Omega$.
@@ VI Comparison with results from finite-element simulations
In Fig. 3 we present streamlines and quiver plots of the leading-order correction to the flow field for a disk undergoing steady rotational motion in a two-dimensional, uniaxially anisotropic fluid. The results are depicted for three values of the parameter $c=\alpha_{\parallel}a$, namely $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). We examine two combinations of the parameters quantifying the anisotropy of viscosity and friction, $A$ and $\beta$, namely $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Perturbative results are displayed in the range $x>0$, while results from finite-element simulations are shown in the range $x<0$.
In the case of vanishing anisotropy of friction, we have demonstrated that $g_{n}$ and $h_{n}$, describing the radial and azimuthal velocities, respectively, vanish for odd $n$. Consequently, $v_{\rho}(\rho,\theta+\pi/2)=v_{\rho}(\rho,\theta)$ and $v_{\theta}(\rho,\theta+\pi/2)=v_{\theta}(\rho,\theta)$, indicating that the velocity remains unchanged upon rotations of $\pi/2$. Anisotropy of viscosity does not notably distort the flow profile with respect to the axis of anisotropy $\bm{\hat{n}}\|\bm{\hat{x}}$, see the top row of Fig. 3. We further observe non-monotonic behavior in the magnitude of velocity, which seems to occur in all situations. This is because the correction $\bm{v}^{*}$ to the flow field vanishes at the circumference of the disk and at infinity. If $\bm{v}^{*}$ does not vanish overall, its magnitude must thus reach a maximum value somewhere in between.
In Fig. 4, we present the leading-order correction to the flow field for rotational motion in a nematic liquid crystal without friction, a scenario for which exact analytical expressions can be derived, see Sec. V. Results are illustrated for the two values of $A=4$ and $10$. They are sufficiently large for the perturbative solution not to remain quantitatively correct any more in general. The analytical solution, displayed on the right-hand sides, aligns perfectly with the finite-element simulations shown on the left-hand sides. Notably, the flow field features regions of large magnitudes of velocity. These regions are arranged along the edge of an octagon near the surface of the disk.
In the listed contexts, we concentrate on presenting the leading-order correction to the flow field due to uniaxial viscosity and friction anisotropy rather than the entire flow field. In this way, we enhance the visualization of the impact of the anisotropies. Our results reveal that the flow field is significantly influenced by changes in the parameters of anisotropy. In some cases, they exhibit vortical structures with closed streamlines and regions of high magnitude of velocity near the disk. Overall, our theoretical predictions based on a perturbative approach align very well with results from finite-element simulations that solve the complete set of hydrodynamic equations.
@@ VII Conclusions
During the further course of the topic and within a wider context, expansions of the theory to thin elastic membranes functionalized by rigid, plate-like inclusions represent a next step [76, 77]. We plan to extend our considerations to elastic composite systems in the future, particularly aiming at a theoretical consideration of membranes of globally aligned liquid crystal elastomers [78, 79, 80, 81, 82, 83].
Finally, in a two-dimensional uniaxially anisotropic fluid film, a disk undergoing rigid rotation would potentially affect the alignment of the axis of anisotropy. We here have considered a simplified and idealized situation, in which the nematic director is assumed to be permanently aligned along a specific direction. This may be due to, for instance, a remotely exerted electric or magnetic field. Our assumption allows the hydrodynamic equations of fluid motion to be considered in a well-addressable form, enabling analytical progress and occasionally exact solutions. Relaxing this assumption to consider a more general scenario, in which the director field is coupled to the velocity field, is a challenging endeavor for the future. As a first step, the situation of a slowly varying director field is particularly interesting, as it may allow for a solution based on perturbative approaches.
@@
The Maple 24 scripts containing the final expressions for the velocity field are accessible through the Zenodo repository using the DOI 10.5281/zenodo.12721219. The data and numerical codes that support the findings of this study are available from the corresponding author upon reasonable request.
Author contribution
A.D.M.I. and A.M.M. designed the research and wrote the manuscript. A.D.M.I. performed the analytical calculations. T.R. conducted the finite-element simulations. A.D.M.I., E.T., M.P., T.R., and A.M.M. reviewed, edited, and provided feedback on the manuscript.